A cylinder of mass m and radius R has a moment of inertia of 1/2mr^2. The cylinder us released from rest at a height h on an inclined plane, and rolls down the plane without slipping. What is the velocity of the cylinder when it reaches the bottom of the incline?

1
Expert's answer
2021-12-09T09:15:15-0500
"mgh=0.5mv^2+0.5I\omega^2\\
mgh=0.5mv^2+0.25mv^2\\v=\sqrt{\frac{4}{3}gh}"